OpenStudy (anonymous):

so when finding the inverse of an equation like x^3+5 do i put the -5 in the square root box with the x?

OpenStudy (anonymous):

\[^3\sqrt{x-5}\] or \[^3\sqrt{x}-5\] and is it \[\frac{ x+6 }{ 7 }\] or is it \[\frac{ x }{ 7 }+6\]

OpenStudy (amistre64):

when do you take the cuberoot? before or after you subtract off the 5?

OpenStudy (amistre64):

you take the cuberoot of a side, not just an element of a side